OpenStudy (anonymous):

Your rich uncle is giving you $100,000 on the condition that you will spend exactly half of the remaining amount each day (or as close to half as possible). Use the equation y = 100,000(0.5)^x to find out how much you will spend on day 8. Round to the nearest cent.

OpenStudy (anonymous):

do you see that x is the number of days? as x increases by 1, the money halfs

OpenStudy (lgbasallote):

substitute x to 8 :D

OpenStudy (anonymous):

go team maths!

OpenStudy (anonymous):

I know but every time I put it into my calculator it says y=0.

OpenStudy (lgbasallote):

lol that's impossible :P a number halved will never reach zero haha

OpenStudy (lgbasallote):

..except zero

OpenStudy (anonymous):

Exactly! Now you understand my confusion :D

OpenStudy (anonymous):

so we do \[100000 \times (0.5)^8 = 390.62500\]

OpenStudy (lgbasallote):

wait..what i said was wrong a while ago..the equation should be y = 100,000(0.5)^7 because the question is how much will you spend ON day 8

OpenStudy (anonymous):

wait, on day 1 we spend 100000*(0.5)

OpenStudy (anonymous):

day 2 we spend 100000*(0.5)^2

OpenStudy (anonymous):

day 3 is 100000*(0.5)^3

OpenStudy (anonymous):

etc, no?

OpenStudy (anonymous):

we cant start with x = 0, we would have no money left lol

OpenStudy (lgbasallote):

good point..ignore what i said

OpenStudy (anonymous):

it was an interesting point